OpenStudy (anonymous):

Will is at a shirt sale where he can buy one and get a second one for 40 percent off. He wants to buy two shirts priced at $29 each. How much will he pay? cost after discount = (2 × list price) – (list price × discount percentage)

OpenStudy (texaschic101):

29 + 0.60(29) = 29 + 17.40 = 46.40

OpenStudy (anonymous):

thank youu

OpenStudy (texaschic101):

its 29 for the first shirt....and he gets 40% off for the second, so he pays 60% for the second....total = 46.40 YW :)